# On the Feasibility of Simple Transformer for Dynamic Graph Modeling

### 25 Jan 2024 v1

@@ 5.5. Hyper-parameter Analysis
We undertake an examination of the critical hyper-parameter choices, taking into account the variations observed across different datasets. Specifically, we systematically explore the impact of several crucial hyper-parameters, namely the number of layers, the number of heads, and the hidden dimension size. These hyper-parameters play a pivotal role in shaping the model’s capacity and its ability to capture intricate patterns within dynamic graphs. We fine-tune the hyper-parameters while keeping all other parameters constant. From Figure 4, we draw some highlights as follows:
@@ 6. Conclusion
In this work, we’ve delved into the intricate realm of dynamic graph modeling, recognizing its profound significance across a range of applications. Drawing from the strengths of the Transformer’s self-attention mechanism, we tailored a solution that sidesteps the often convoluted designs prevalent in existing methods. Our novel approach re-envisions dynamic graphs from a sequence modeling perspective, leading to the development of an innovative temporal alignment technique. This strategic design not only adeptly captures the temporal dynamics inherent in evolving graphs but also simplifies their modeling process. Our empirical investigations, carried out across four real-world datasets spanning diverse sectors, serve as a testament to our model’s efficacy. In the future, we will delve deeper into the nuances of the temporal alignment technique for further optimizations. Additionally, the potential for integrating more advanced attention mechanisms can be explored to further elevate the model’s capabilities in capturing dynamic evolutions.
Appendix A Additional Implement Details
@@
Appendix B Hyper-parameters Settings of Baselines

### 27 Feb 2024 v2

@@ 5.5. Hyper-parameter Analysis
We undertake an examination of the critical hyperparameter choices, taking into account the variations observed across different datasets. Specifically, we systematically explore the impact of several key hyperparameters, namely, the number of layers, the number of heads, and the hidden dimension size. These hyperparameters play a pivotal role in shaping the model’s capacity and its ability to capture intricate patterns within dynamic graphs. We vary the value of each hyperparameters while keeping all other parameters constant. From Figure 4, we draw some key observations as follows.
@@ 6. Conclusion
In this work, we explored the problem of dynamic graph modeling, recognizing its significance across a wide range of applications. Drawing from the strengths of the Transformer’s self-attention mechanism, we tailored a solution that supersedes the often convoluted designs in many existing methods. Our novel approach, named SimpleDyG, reformulates dynamic graphs from a sequence modeling perspective. It is superior to not only the discrete-time approaches by considering the full temporal order, but also the continuous-time approaches by capturing long-term dependencies using a Transformer. SimpleDyG is exceptionally simple as it does not modify the Transformer architecture; instead, it maps a dynamic graph into a set of sequences via temporal ego-graphs, and modifies the input sequences to achieve temporal alignment. Nevertheless, SimpleDyG achieves surprisingly strong performance in diverse dynamic graphs. As future work, we will investigate the nuances of the temporal alignment technique for further optimizations.
Acknowledgments
This research / project is supported by the Ministry of Education, Singapore, under its Academic Research Fund Tier 2 (Proposal ID: T2EP20122-0041). Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the views of the Ministry of Education, Singapore.
Appendices
Appendix A Link Deletion Operation
| Method | DySAT | EvolveGCN | DyRep | JODIE | TGAT | TGN | TREND | GraphMixer | SimpleDyG |
| Time (s) | 12.24 | 11.89 | 6.4 | 6.25 | 18.54 | 8.05 | 7.45 | 6.89 | 6.21 |
Our framework can support the link deletion operation within the evolution of dynamic graphs. For example in the ML-10M dataset with links representing users’ ratings for the movies, assuming that users can delete the historical ratings. Suppose the historically rated movies of one user is “”, he/she may delete the rating of b, then we can add a special token “[del]” to identify the deletion operation: “”. It is worth noting that most works for dynamic graph modeling are associated with link addition with benchmark datasets mainly involving this type of operation.
Appendix B Additional Implement Details
@@
Appendix C Hyperparameters Settings of Baselines
@@
Appendix D Statistical Temporal Pattern Analysis in Each Dataset
We incorporate statistical analyses to understand the temporal patterns within each dataset. Specifically, we counted the number of interactions per day for the UCI, ML-10M, and Hepth datasets. For the conversation dataset (MMConv), we counted the number of interactions per turn. We split the dataset into several bins with each bin covering a range of time (10 days, 90 days, 60 days and two turns for UCI, ML-10M, Hepth and MMConv, respectively), and then calculate the average daily interaction counts in each bin. We show the results in Figure A.1.
We observe sudden changes in interactions in the UCI and ML-10M datasets. While the Hepth and MMConv datasets generally show gradual change across the entire timeline. This indicates that in the UCI and ML-10M datasets, the interactions among different time slots show different patterns. Thus it is more important to conduct temporal alignment using temporal tokens which is consistent with our analysis for Table 4. For the Hepth and MMConv datasets, a simpler design for temporal alignment (e.g., same token or no token) is enough to achieve good performance.
Appendix E Time Complexity Analysis
The time complexity of our SimpleDyG model is the same as the vanilla Transformer, which is where is the sequence length. We conduct time efficiency experiments about the training time per epoch of different methods on UCI dataset using a machine with a NVIDIA L40 GPU with 64 CPU cores. The results show that our method trains faster than or comparable to all baselines. The methods integrating temporal modeling (e.g., RNN, self-attention) with structural modeling (e.g., GNN, GAT) suffer from the complexity of these modules.